Sue and Carol start bicycling toward each other at the same time from homes 9 miles apart. When they meet, Carol has traveled c miles, and Sue has traveled __________ miles.
A. 9
B. c – 9
C. 9 – c
D. 9 + c
Mary Beth and Michael leave school traveling in opposite directions. Michael is walking and Mary Beth is biking, averaging 6 km/h more than Michael. If they are 18 km apart after 1.5 hours, what is Michael's rate?
A. 9 km/h
B. 3 hours
C. 9 mi/h
D. 3 km/h
Juan left home on his bicycle at 10:00 a.m. traveling at 21 km/h. At noon, his brother set out after him on his motorcycle following the same route. If the motorcycle traveled at 63 km/h, what time did Juan's brother overtake him?
A. 12:30 p.m.
B. 11:00 a.m.
C. 1:00 p.m.
D. 1:00 a.m.
Thank you so much, sir !!!

Answer by stanbon(75887)

Sue and Carol start bicycling toward each other at the same time from homes 9 miles apart. When they meet, Carol has traveled c miles, and Sue has traveled
9-c miles.
A. 9
B. c – 9
C. 9 – c
D. 9 + c
----------------------
Mary Beth and Michael leave school traveling in opposite directions. Michael is walking and Mary Beth is biking, averaging 6 km/h more than Michael. If they are 18 km apart after 1.5 hours, what is Michael's rate?
---
Michael rate = x km/h ; Mary's rate = (x+6) km/h
Michael distance: 1.5x km
Mary distance::::1.5(x+6)
Equation:
1.5x + 1.5(x+6) = 18 km
1.5x + 1.5x + 9 = 18
3x = 9
x = 3 km/hr (Michael's rate)
===============================
A. 9 km/h
B. 3 hours
C. 9 mi/h
D. 3 km/h
=====================================
Juan left home on his bicycle at 10:00 a.m. traveling at 21 km/h. At noon, his brother set out after him on his motorcycle following the same route. If the motorcycle traveled at 63 km/h, what time did Juan's brother overtake him?
---
Juan rate = 21 km/h ; time = x hrs ; distance = rt = 21x km
---
Brother rate = 63 km/h ; time = x-2 hrs ; distance = 63x - 126 km
------------------------------
Equation:
distance = distance
21x = 63x - 126
42x = 126
x = 3 hrs
--------
Overtake time:
10:00a.m. + 3 hrs = 1:00p.m.
===============================
Cheers,
Stan H.

A. 12:30 p.m.
B. 11:00 a.m.
C. 1:00 p.m.
D. 1:00 a.m.
=================================
Cheers,
Stan H.